# NBBOX: Noisy Bounding Box Improves Remote Sensing Object Detection

### v1

@@ III-D Comparison with state-of-the-art methods
Data DIOR-R DOTA
Detectors Faster R-CNN FCOS RetinaNet Faster R-CNN
Baseline 62.66 ± 0.17 59.75 ± 1.09 55.69 ± 0.77 68.92 ± 0.44
+ RandRotate [30] 63.10 ± 0.21 60.11 ± 1.12 55.85 ± 0.87 70.02 ± 0.32
+ RandShift [30] 62.87 ± 0.37 59.67 ± 1.05 55.07 ± 0.83 69.01 ± 0.40
+ Cutout [3] 62.91 ± 0.26 59.84 ± 0.99 55.83 ± 0.92 69.13 ± 0.45
+ AutoAug [31] 62.33 ± 0.31 59.24 ± 1.07 54.78 ± 1.13 68.87 ± 0.32
+ RandAug [32] 62.55 ± 0.12 59.94 ± 1.29 55.17 ± 1.22 66.11 ± 0.24
+ AugMix [33] 62.83 ± 0.11 59.93 ± 0.75 55.74 ± 0.99 69.26 ± 0.10
+ MosaicMix [6, 34] 61.15 ± 0.19 58.03 ± 0.76 56.15 ± 0.31 62.73 ± 0.58
+ YOCO [35] 57.52 ± 0.16 54.37 ± 1.70 51.14 ± 1.01 65.16 ± 0.36
+ PixMix [36] 62.07 ± 0.23 58.91 ± 1.13 54.61 ± 0.74 68.72 ± 0.33
+ Ours (Best) 63.18 ± 0.14 60.25 ± 1.00 56.09 ± 0.61 69.79 ± 0.35
+ Ours (Best) + RandRotate - - - 70.36 ± 0.31
+ Ours (Best) + MosaicMix - - 56.32 ± 0.36 -
TABLE V: Time analysis of each transformation used in this work.
Detector: Faster R-CNN DIOR-R DOTA
RandRotate [30] 09.65min 17.12min
RandShift [30] 09.66min 16.37min
Cutout [3] 09.66min 16.29min
AutoAug [31] 09.70min 16.32min
RandAug [32] 09.64min 16.34min
AugMix [33] 09.75min 16.61min
MosaicMix [6, 34] 21.24min 34.48min
YOCO [35] 09.77min 16.40min
PixMix [36] 09.75min 16.54min
Ours (Scaling + Rotation + Translation) 09.63min 16.28min
@@ IV Conclusion
In this letter, we have proposed a novel data augmentation method named NBBOX for remote sensing object detection. This work thoroughly investigates the effect of bounding box transformation with scaling, rotation, and translation for remote sensing object detection. Extensive experiments on DOTA and DIOR-R demonstrate that noise injection into bounding boxes surpasses existing state-of-the-art augmentation strategies based on image manipulation. Besides, NBBOX is easy to integrate with modern deep-learning frameworks and more time-efficient than other augmentation methods. In the future, we will extend our method with self-supervised learning such as multiple instance learning for further improvement.

### v3

@@ III-E Comparison with state-of-the-art methods
Backbone Data DIOR-R DOTA
Detectors Faster R-CNN FCOS Faster R-CNN FCOS
ResNet-50 Baseline 62.66 ± 0.17 59.75 ± 1.09 68.92 ± 0.44 67.00 ± 0.26
+ RandRotate [33] 63.10 ± 0.21 60.11 ± 1.12 70.02 ± 0.32 67.56 ± 0.49
+ RandShift [33] 62.87 ± 0.37 59.67 ± 1.05 69.01 ± 0.40 67.22 ± 0.59
+ CutOut [5] 62.91 ± 0.26 59.84 ± 0.99 69.13 ± 0.45 67.42 ± 0.37
+ AutoAug [34] 62.33 ± 0.31 59.24 ± 1.07 68.87 ± 0.32 67.45 ± 0.49
+ RandAug [35] 62.55 ± 0.12 59.94 ± 1.29 66.11 ± 0.24 65.08 ± 0.70
+ AugMix [36] 62.83 ± 0.11 59.93 ± 0.75 69.26 ± 0.10 66.71 ± 0.52
+ MosaicMix [8, 37] 61.15 ± 0.19 58.03 ± 0.76 62.73 ± 0.58 61.18 ± 0.57
+ UniformAug [38] 59.18 ± 0.18 56.35 ± 1.53 65.61 ± 0.37 64.35 ± 0.84
+ TrivialAug [39] 60.22 ± 0.17 57.37 ± 1.48 66.54 ± 0.19 65.34 ± 0.56
+ YOCO [40] 57.52 ± 0.16 54.37 ± 1.70 65.16 ± 0.36 63.12 ± 0.48
+ PixMix [41] 62.07 ± 0.23 58.91 ± 1.13 68.72 ± 0.33 66.82 ± 0.28
+ InterAug [42] 60.94 ± 0.27 58.45 ± 1.20 66.93 ± 0.24 66.52 ± 0.71
+ Ours 63.21 ± 0.18 60.27 ± 1.06 69.78 ± 0.54 67.89 ± 0.46
+ Ours + RandRotate - - 70.37 ± 0.13 -
Swin-T Baseline 61.53 ± 0.49 63.60 ± 0.99 71.03 ± 0.47 68.40 ± 0.68
+ RandRotate [33] 62.54 ± 0.43 64.08 ± 0.78 71.62 ± 0.34 68.52 ± 0.95
+ RandShift [33] 62.77 ± 0.30 63.22 ± 0.73 71.21 ± 0.50 68.51 ± 0.64
+ Cutout [5] 63.17 ± 0.36 63.80 ± 0.88 71.29 ± 0.28 68.41 ± 0.82
+ AutoAug [34] 62.80 ± 0.43 63.19 ± 0.94 70.78 ± 0.64 69.04 ± 0.72
+ RandAug [35] 60.54 ± 0.60 62.07 ± 0.95 68.63 ± 0.41 68.91 ± 0.68
+ AugMix [36] 62.55 ± 0.27 63.07 ± 1.01 70.48 ± 0.61 69.06 ± 0.57
+ MosaicMix [8, 37] 58.26 ± 0.31 60.57 ± 0.76 66.48 ± 0.39 67.24 ± 0.72
+ UniformAug [38] 58.48 ± 0.21 60.55 ± 0.91 67.38 ± 0.45 69.06 ± 0.71
+ TrivialAug [39] 59.72 ± 0.27 61.44 ± 1.12 67.63 ± 0.42 68.53 ± 0.69
+ YOCO [40] 57.05 ± 0.32 58.79 ± 1.09 66.26 ± 0.39 69.04 ± 0.59
+ PixMix [41] 58.15 ± 0.31 58.22 ± 0.81 66.90 ± 0.60 68.15 ± 0.32
+ InterAug [42] 60.72 ± 0.20 62.24 ± 0.99 69.35 ± 0.38 67.09 ± 0.84
+ Ours 62.22 ± 0.23 63.64 ± 0.89 71.30 ± 0.35 68.94 ± 0.59
+ Ours + Cutout 63.18 ± 0.30 - - -
+ Ours + RandRotate - 64.11 ± 0.98 71.69 ± 0.41 -
+ Ours + AugMix - - - 69.12 ± 0.50
@@ IV Conclusion
In this letter, we have proposed a novel data augmentation method named NBBOX for remote sensing object detection. This work thoroughly investigates the effect of bounding box transformation with scaling, rotation, and translation for remote sensing object detection. Besides, NBBOX is easy to integrate with modern deep-learning frameworks and more time-efficient than other augmentation methods. In the future, we will extend our method with various AI paradigms.